Exactly how were the MPC and MPS in Table 21-4 computed? Illustrate by calculating MPC and MPS between points A and B. Explain why it must always be true that MPC  MPS  1 .


1
Expert's answer
2021-09-27T18:40:25-0400

There is an increase in savings when income increases. The consumption can also increase as a result of increase in income.

Marginal Propensity to Consume (MPC)

"MPC = \frac{Consumption \ Change}{Income\ Change}"

"Consumption \ at \ Point \ A = 24200"

"Consumption \ at \ Point \ B = 25000"

"Consumption \ Change= 25000 - 24200"

"Consumption \ Change= 800"

"Point \ A \ income = 24000"

"Point \ B \ income = 25000"

"Income\ change = 25000 - 24000 = 1000"

"MPC = \frac{800}{1000}"

"MPC =0.8"

Marginal Propensity to Save (MPS)

"MPS = \frac{Change\ in \ Savings}{Income\ Change}"

"Savings \ at \ Point \ A = -200 \\ Savings\ at \ Point = 0 \\ Change \ in \ savings = 0 -- 200 = 200
\\ Income \ at \ Point\ A = 24000 \\ Income \ at\ point \ B = 25000
\\ Change \ in \ Income = 25000 - 24000 = 1000 \\ 
MPS= \frac{200}{1000} = 0.2 \\ 0.2 + 0.8 = 1 \\ Therefore, it\ is\ true\ that\ MPC+ MPS = 1, \ since \\
0.8 + 0.2 = 1"